Solve the following equation so that it can be plotted on a graphing calculator.
4x^2 - 2xy + 3y^2 =2. I know it should be able able to use the quadratic formula but every way I try, I do not get the right answer. The question came from the TeXeS Review manual. There is no ISBN.

Answer by AnlytcPhil(1806)

Solve the following equation so that it can be plotted on a graphing calculator.
4x^2 - 2xy + 3y^2 =2. I know it should be able able to use the quadratic
formula but every way I try, I do not get the right answer. The question came
from the TeXeS Review manual. There is no ISBN.

4x² - 2xy + 3y² = 2

Rearrange it in descending order of powers of y with 0 on the right:

3y² - 2xy + 4x² - 2 = 0

Write it this way, to see what A, B, and C are.

(3)y² + (-2x)y + (4x²-2) = 0

Then for the quadratic formula,

A = (3),  B = (-2x),  C = (4x²-2)
           ________
     -B ± ÖB² - 4AC
y = —————————————————
           2A

               ____________________
     -(-2x) ± Ö(-2x)² - 4(3)(4x²-2)
y = —————————————————————————————————
                2(3)

You can put that in the TI-83 just like this without simplifying further:

Y1 = (-(-2X) + Ö((-2X)² - 4(3)(4X²-2)))/(2(3))

Y2 = (-(-2X) - Ö((-2X)² - 4(3)(4X²-2)))/(2(3))

use window Xmin = -1, Xmax = 1, Xscl = 1, Ymin = -1, Ymax = 1, Yscl=1, Xres=1

You get a slanted ellipse

or you can simplify it further first

               ____________________
     -(-2x) ± Ö(-2x)² - 4(3)(4x²-2)
y = —————————————————————————————————
                2(3)

           _______________
     2x ± Ö4x² - 12(4x²-2)
y = ———————————————————————
                6

           _______________
     2x ± Ö4x² - 48x² + 24
y = ———————————————————————
                6

           ___________
     2x ± Ö-44x² + 24
y = ———————————————————
             6

           ___________
     2x ± Ö4(-11x² + 6
y = ———————————————————
             6

            ___________
     2x ± 2Ö(-11x² + 6
y = ———————————————————
             6

            __________
     2[x ± Ö-11x² + 6
y = ———————————————————
             6

Cancel the 2 into the 6

          _________
     x ± Ö-11x² + 6
y = —————————————————
            3

          ________
     x ± Ö6 - 11x² 
y = ———————————————
            3

Then you can enter it in the TI-83 as
            
Y1 = (X + Ö(6 - 11X²))/3

Y2 = (X - Ö(6 - 11X²))/3

Either way you get this slanted ellipse.



Edwin